OpenStudy (anonymous):

add, simplify if possible 4w/w^2-9 + w/w-3

jimthompson5910 (jim_thompson5910):

\[\Large \frac{4w}{w^2-9} + \frac{w}{w-3}\] \[\Large \frac{4w}{(w-3)(w+3)} + \frac{w}{w-3}\] \[\Large \frac{4w}{(w-3)(w+3)} + \frac{w(w+3)}{(w-3)(w+3)}\] \[\Large \frac{4w}{(w-3)(w+3)} + \frac{w^2+3w}{(w-3)(w+3)}\] \[\Large \frac{4w+w^2+3w}{(w-3)(w+3)}\] \[\Large \frac{w^2+7w}{(w-3)(w+3)}\] \[\Large \frac{w^2+7w}{w^2-9}\] So \[\Large \frac{4w}{w^2-9} + \frac{w}{w-3}=\frac{w^2+7w}{w^2-9}\]